using weierstrass m-test show that the following series converges uniformly∑ n=1 ∞ n^3 X^n,X€[-1/3,1/3]


1
Expert's answer
2021-06-11T03:07:55-0400

Given,"\sum_{n=1}^{\infty} n^3x^n, x\in[-\frac{1}{3},\frac{1}{3}].\\
Then,\\
f_n=n^3x^n\to\infty as n\to \infty\\
i.e., \nexists M_n \\
\text{such that,}\\
f_n(x)|\leq M_n, \forall n\in \N, \forall x\in[-\frac{1}{3},\frac{1}{3}].\\
\text{Thus, by weierstrass m-test the given series is not uniformly covergent.}\\"